OpenStudy (anonymous):

Algebra 1 Help! Will Fan and Medal!

OpenStudy (anonymous):

There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B. Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points) Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points) Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. (4 points)

OpenStudy (lukeskylicker):

First off I spot test question.

OpenStudy (janu16):

Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points) y = 30 (1.2)^x y= 45 + 3*x Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points) 30(1.2)^5= 74.65 45 + 3*5= 60 Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. solve 30 (1.2)^x = 45 + 3*x using a graphing calculator i got x = 3.32079 in the third year the number of homes are the same

OpenStudy (lukeskylicker):

Second of all I will explain to you how to get the answers for Part A through C but not any actual answers. Part A: y will equal the number of homes so we have y = that will be the beginning of both answers. Afterwards there will be the constant (the total number of homes when x is equal to zero) and the amount increasing the total when x is equal to a positive number. Remember though. If the equation is increasing linearly do x (the year) times the amount number your given (or put another way, if your given an integer use multiplication) but if it's increasing exponentially have the number (which should be a percentage) to the power of x or \[y^x\] Part B: take the equation from Part A and replace all instances of x with 5. Part C: input your functions into a graphing calculator or use desmos.com

OpenStudy (anonymous):

Thank you.